# WMCopier: Forging Invisible Image Watermarks on Arbitrary Images

### v2

@@ D.2 Discrimination of Forged Watermarks by Robustness Gap
Appendix E Potential Defenses
E.1 Semantic Watermark
Semantic watermarking [46, 48] embeds watermark information that is intrinsically tied to the semantic content of the image. This tight coupling makes it substantially more difficult to forge the watermark onto arbitrary images. Table 6 presents the results of our WMCopier on the semantic watermark TreeRing [46].
Table 6: Attack performance on TreeRing [46].
@@ E.1 Semantic Watermark
MS-COCO 33.60 0.00% 5.00%
CelebA-HQ 31.65 0.00% 6.00%
ImageNet 33.62 0.00% 3.00%
DiffusionDB 33.46 0.00% 6.00%
The results suggest that our attack fails to forge the TreeRing watermark. This highlights the potential of semantic watermarking as a defense against our attack. Although semantic watermarking is still in the early stages of research and has yet to see widespread practical deployment, we encourage future work to further explore and advance this promising direction.
E.2 Attack on Multiple Messages
To enhance the deployed watermarking system, we suggest modifying the existing watermark system by disrupting the ability of diffusion models to model the watermark distribution effectively. Specifically, we propose a multi-message strategy as a simple yet effective countermeasure. Instead of embedding a fixed watermark message, the system randomly selects one from a predefined message pool $m_{1},m_{2},m_{3},\dots,m_{K}$ for each image. During detection, the detector verifies the presence of any valid message in the pool. This strategy introduces uncertainty into the watermark signal, increasing the entropy of possible watermark patterns and making it substantially more difficult for generative models to learn consistent features necessary for forgery. We implement this defense using different message pool sizes ($K=10,50,100$).
Table 7: Performance comparison across different $K$ values.
K=10 K=50 K=100
Dataset PSNR↑ Forged Bit-acc.↑ FPR@0.05↑ PSNR↑ Forged Bit-acc.↑ FPR@0.05↑ PSNR↑ Forged Bit-acc.↑ FPR@0.05↑
MS-COCO 34.74 80.14% 88.90% 32.85 69.89% 74.80% 34.97 69.62% 72.70%
CelebAHQ 36.10 83.75% 96.70% 33.71 70.43% 80.20% 35.97 70.17% 77.50%
ImageNet 34.67 79.24% 87.20% 32.78 69.82% 77.20% 34.87 69.60% 72.70%
Diffusiondb 35.21 76.17% 79.40% 33.32 70.09% 76.20% 35.74 69.73% 74.70%
Although increasing the value of $K$ leads to a notable reduction in FPR, we further enhance our attack by collecting more watermarked images. Specifically, we gather 5,000, 20,000, and 50,000 watermarked samples to evaluate the impact of data volume on attack performance. As shown in Table 8, the attack remains highly effective when using a large auxiliary set $\mathcal{D}_{aux}$, achieving over 90% success rate. While this highlights the vulnerability of the watermarking scheme, it also suggests a potential weak defense mechanism: increasing the cost of the attack by requiring a larger number of collected watermarked images (e.g., generating 20,000 images may cost around \$800).
Dataset 5000 20000 50000
PSNR↑ Forged Bit-acc.↑ FPR@0.05↑ PSNR↑ Forged Bit-acc.↑ FPR@0.05↑ PSNR↑ Forged Bit-acc.↑ FPR@0.05↑
MS-COCO 34.97 69.62% 72.70% 33.27 72.19% 93.20% 30.86 72.24% 95.00%
CelebA-HQ 35.97 70.17% 77.50% 30.87 72.65% 96.00% 29.57 72.48% 94.50%
ImageNet 34.87 69.60% 72.70% 33.21 72.04% 92.90% 30.68 72.21% 93.50%
DiffusionDB 35.74 69.73% 74.70% 33.21 72.06% 93.00% 31.08 72.46% 93.90%
Appendix F Ablation Study
Table 9 shows that the refinement significantly improves visual fidelity (PSNR) and forgery effectiveness (forged bit accuracy).
Table 9: Impact of refinement on forgery performance.
Watermark Scheme PSNR Forged Bit-acc. FPR@0.05
@@ Appendix F Ablation Study
DWT-DCT 32.40 33.77 63.03% 89.62% 39.00% 95.00%
HiddeN 29.81 32.79 80.60% 99.40% 92.00% 100.00%
RivaGAN 31.89 34.03 89.90% 95.90% 95.00% 97.00%
StableSignature 25.60 31.27 97.58% 98.19% 100.00% 100.00%
Appendix G Training Details of the Diffusion Model
We adopt a standard DDIM framework for training, following the official Hugging Face tutorial [^3] HuggingFace Tutorial: https://huggingface.co/docs/diffusers/en/tutorials/basic_training. The model is trained for 20,000 iterations with a batch size of 256 and a learning rate of $1\times 10^{-4}$. To support different watermarking schemes, we only adjust the input resolution of the model to match the input dimensions for each watermark. Other training settings and model configurations remain unchanged. Although the current training setup suffices for watermark forgery, enhancing the model’s ability to better capture the watermark signal is left for future work. For our primary experiments, we train an unconditional diffusion model from scratch using 5,000 watermarked images. Due to the limited amount of training data, the diffusion model demonstrates memorization [6], resulting in reduced sample diversity, as illustrated in Figure 8. All of the experiments are conducted on an NVIDIA A100 GPU.
Appendix H Limitation
In this section, we discuss the limitations of WMCopier. While our current training paradigm already achieves effective watermark forgery, we have yet to systematically investigate how to better guide diffusion models to capture the underlying watermark distribution. Moreover, although our method is effective against widely adopted invisible watermarking schemes, it may be less effective against semantic watermarking strategies, which embed information into high-level image semantics. While such techniques are not yet widely deployed in practice, they present a promising direction for future defense.
Appendix I Broader Impact
@@
Appendix J Forged Samples

### v3

@@ D.3 Discrimination of Forged Watermarks by Robustness Gap
Appendix E Additional Ablation Studies
Table 11 shows that the proposed refinement step substantially improves visual fidelity, as measured by PSNR, while simultaneously enhancing forgery performance (forged-bit accuracy).
We also explore the impact of varying the size of $D_{\mathrm{aux}}$. Specifically, we use 1,0001{,}000, 5,0005{,}000, and 10,00010{,}000 collected RivaGAN watermarked images. As shown in Table 12, larger $D_{\mathrm{aux}}$ generally yields higher forged-bit accuracy and higher FPR across datasets. However, the improvement becomes marginal once the size of $D_{\mathrm{aux}}$ reaches around 5,0005{,}000, indicating that the attack performance saturates beyond this point.
Table 11: Impact of refinement on forgery performance.
@@ Appendix E Additional Ablation Studies
DWT-DCT 32.40 33.77 63.03% 89.62% 16.00% 57.00%
HiddeN 29.81 32.79 80.60% 99.40% 89.00% 94.00%
RivaGAN 31.89 34.03 89.90% 95.90% 84.00% 96.00%
StableSignature 25.60 31.27 97.58% 98.19% 91.00% 98.00%
Table 12: Performance comparison across datasets with different sizes of $D_{aux}$
Dataset 1000 5000 10000
[…]
MS-COCO 34.16 81.82% 80.70% 34.07 95.74% 96.40% 34.47 97.81% 96.30%
CelebA-HQ 35.74 89.10% 89.50% 35.28 98.61% 99.10% 35.25 98.63% 98.50%
ImageNet 34.10 81.25% 71.50% 33.87 93.83% 94.90% 34.29 93.53% 95.80%
DiffusionDB 34.77 74.76% 64.10% 34.50 90.43% 91.20% 34.96 91.70% 93.60%
Appendix F Training Details of the Diffusion Model
We adopt a standard DDIM framework for training, following the official Hugging Face tutorial [^3] HuggingFace Tutorial: https://huggingface.co/docs/diffusers/en/tutorials/basic_training. The model is trained for 20,000 iterations with a batch size of 256 and a learning rate of $1\times 10^{-4}$. The entire training process takes roughly 40 A100 GPU hours. To support different watermarking schemes, we only adjust the input resolution of the model to match the input dimensions for each watermark. Other training settings and model configurations remain unchanged. Although the current training setup suffices for watermark forgery, enhancing the model’s ability to better capture the watermark signal is left for future work. For our primary experiments, we train an unconditional diffusion model from scratch using 5,000 watermarked images. Due to the limited amount of training data, the diffusion model demonstrates memorization [18], resulting in reduced sample diversity, as illustrated in Figure 9. All of the experiments are conducted on an NVIDIA A100 GPU.
Appendix G Limitation
In this section, we discuss the limitations of our attack. While our current training paradigm already achieves effective watermark forgery, we have not yet systematically explored how to guide diffusion models better to capture the underlying watermark distribution. In this work, we employ a standard diffusion architecture without any specialized training strategies. We leave the exploration of alternative architectures and training schemes to future work. Moreover, understanding why different watermark types exhibit varying forgery and learning behaviors remains an open problem. Additionally, our method requires a substantial amount of data and incurs training costs.
Appendix H Broader Impact
@@
Appendix I Forged Samples